OpenStudy (anonymous):

If Zack makes 1½ quarts of lemonade, then he uses 6 lemons. If Zack makes 1½ quarts of lemonade, then he makes 4 servings. Zack uses 5 lemons. Which is a valid conclusion? a.Zack does not make 1½ quarts. b.Zack makes 3 servings. c.Zack makes 1 quart. d.Zack makes 2 servings.

OpenStudy (anonymous):

"A" is a possibility "B" and "D" have to do with servings.. 2 servings would take 1/2 the number of lemons which is 1/2 * 6 = 3 lemons. 3 servings would take 3/4 the number of lemons which would be 3/4 of 6 which is 3/4 * 6 = 18/4 = 4.5 lemons. "C" making a quart would be 2/3 of 6 = 4 lemons. NOW.. this is your choice. A is obviously wrong, but if you think that if you needed 4.5 lemons you would put in 5 lemons, then B would be your answer.